OpenStudy (anonymous):

Consider the function y(x) = 5 cos(x/3+π/4)  . a) Find the period. 1. None of these 2.6/π 3. 3π 4. 5 5.3/π 6. 6

OpenStudy (anonymous):

period is \[\frac{2\pi}{b}=\frac{2\pi}{\frac{1}{3}}=6\pi\]

OpenStudy (anonymous):

so looks like none of the above. is this some on line class?

OpenStudy (turingtest):

yes, the answer is definitely \[6\pi\]so there is an error either on the test or your post.